# Indecomposable set-theoretical solutions to the Yang–Baxter equation of size $p^{2}$

### v1

@@ 4. Indecomposable irretractable cycle sets of size $p^{2}$ The general case
From the original definition of $\operatorname{Fix}_{A}(L\cap G)$ we see that $\operatorname{Fix}_{A}(L\cap G)^{+}$ is a group. Let $a\in\operatorname{Fix}_{A}(L\cap G)$ and $b\in L\cap G$. Then $\lambda_{\overline{a}}(b)\in L$, but as $\lambda_{\overline{a}}(b)={}^{\overline{a}}b$, also $\lambda_{\overline{a}}(b)\in G$. We conclude that $\lambda_{\overline{a}}(b)\in L\cap G$. Using the alternative description of $\operatorname{Fix}_{A}(L\cap G)$ we now see that $\operatorname{Fix}_{A}(L\cap G)^{\circ}$ is multiplicatively closed and non-empty. As $A$ is finite, we conclude that $\operatorname{Fix}_{A}(L\cap G)^{\circ}$ is a subgroup of $A^{\circ}$. Therefore $\operatorname{Fix}_{A}(L\cap G)$ is a subbrace. As ${}^{\overline{a}}b\in L\cap G$ for all $a\in\operatorname{Fix}_{A}(L\cap G)$, $b\in L\cap G$ we find indeed that $\operatorname{Fix}_{A}(L\cap G)$ is contained in the normaliser of $L\cap G$ in $A^{\circ}$. ∎
In the remainder of the section, we let $X$ be an indecomposable irretractable cycle set of size $p^{2}$ and $\mathcal{G}=\mathcal{G}(X)$. As we already covered the case where $\mathcal{G}$ is a $p$-group in Section 3, we can assume that $\mathcal{G}$ is not a $p$-group. We associate $X$ with its image in $\mathcal{G}$, which is a transitive cycle base. As $X$ is irretractable, it follows from [2, Lemma 2.1.] that $\operatorname{Soc}(\mathcal{G})=0$.
@@
From now on we consider the unique block system of $X$ under the action of $\mathcal{G}$, and recall that the uniqueness is guaranteed by 1.7. As before we denote this by $\{\mathcal{B}_{x}\mid x\in X\}$. We denote the subgroup of $\mathcal{G}^{\circ}$ that fixes the blocks setwise by $\mathcal{A}$. Also, we define $\mathcal{A}_{p}=\mathcal{A}\cap\mathcal{G}_{p}$ and $\mathcal{A}_{p^{\prime}}=\mathcal{A}\cap\mathcal{G}_{p^{\prime}}$. Note that in particular, $\mathcal{A}_{p}$ is a $p$-Sylow subgroup of $\mathcal{A}^{\circ}$ and $\mathcal{A}_{p^{\prime}}$ is a Hall $p^{\prime}$-subgroup of $\mathcal{A}^{\circ}$. In particular, $\mathcal{A}_{p}$ is normal in $\mathcal{G}^{\circ}$ by 1.8.
@@
In particular, from 1.8 we find that $|\mathcal{G}_{p^{\prime}}|<p$ and thus the $\lambda$-action of $\mathcal{G}_{p}$ on $\mathcal{G}_{p^{\prime}}$ is trivial. As this same action is transitive on $Z$ we find $|Z|=1$. Let $z\in Z$, then $\lambda_{z}$ is a brace automorphism of $\mathcal{G}_{p}$ by 1.1 and therefore also its restriction to $Y$ yields a cycle set automorphism. For any $x_{i},x_{j}\in X$ we find
@@
We remark that the idea of starting from a cycle set $X$ and then considering its projection $Y$ onto the $p$-Sylow subgroup $\mathcal{G}_{p}$, is strongly related to the notion of cabling, see [15, 10]. Recall that for $k\in\mathbb{Z}$, the $k$-cabled cycle set is defined as $(X,\cdot_{k})$ with $x\cdot_{k}y=(k\sigma_{x})(y)$, where $k\sigma_{x}\in\mathcal{G}(X)$ is the $k$-th power of $\sigma_{x}$ in $(\mathcal{G}(X),+)$.
@@
Let $|\mathcal{G}(X,\cdot)|=p^{r}m$ where $(p,m)=1$. As the additive $p$-Sylow subgroup is $\mathcal{G}(X,\cdot)_{p}=\{kg\mid g\in\mathcal{G}\}$ for any non-zero multiple $k$ of $m$, we find that $\mathcal{G}(X,\cdot)_{p}=\mathcal{G}(X,\cdot_{k})$. In particular, $\operatorname{Ret}(X,\cdot_{k})$ isomorphic to the cycle set $\{k\sigma_{x}\mid x\in X\}\subseteq\mathcal{G}(X,\cdot)$ where the equivalence class of $x$ is mapped to $k\sigma_{x}$. If we let $k$ be such that $m\equiv 1\pmod{p^{r}}$ and $m\equiv 0\pmod{k}$ then we find that the $k\sigma_{x}$ is precisely the projection of $\sigma_{x}$ onto $\mathcal{G}(X,\cdot)_{p}$, hence in this case we find that $Y\cong\operatorname{Ret}(X,\cdot_{k})$.
@@
By the functoriality of $G(X,r)$ we get an induced automorphism $\phi^{\prime}$ of $G(X,r)$ which restricts to $\phi$ on the generating set $X\subseteq G(X,\cdot)$. We let $\mathbb{Z}$ denote the trivial brace on $\mathbb{Z}$ and we let $\mathbb{Z}^{\circ}$ act on $G(X,\cdot)$ by $1\cdot g=(\phi^{\prime})^{-1}(g)$ for $g\in G(X,\cdot)$. Hence we can construct the semi-direct product $G(X,\cdot)\rtimes\mathbb{Z}$ in the sense of [19, Corollary 3.37]. The set $X+1\subseteq G(X,\cdot)\rtimes\mathbb{Z}$ is closed under the $\lambda$-action. Hence it is a sub-cycle set which is precisely the cycle set in the statement under the correspondence $x\mapsto x+1$.
@@
Let $(X,\cdot)$ be a cycle set and $\phi$ an automorphism of $(X,\cdot)$ of order $m$ coprime to $|\mathcal{G}(X)|$ such that $\phi$ has a fixed point. Then $\mathcal{G}(X,\cdot_{\phi})^{\circ}=\mathcal{G}(X,\cdot)^{\circ}\rtimes\langle\phi\rangle$ as subgroups of $\operatorname{Sym}_{X}$ and $\mathcal{G}(X,\cdot_{\phi})\cong\mathcal{G}(X,\cdot)\rtimes\mathbb{Z}_{m}$ as braces where $\mathbb{Z}_{m}$ is given the trivial brace structure and acts on $\mathcal{G}(X,\cdot)$ by $\phi$.
@@
Let $(X,\cdot)$, $(X,\cdot^{\prime})$ be cycle sets and $\phi\in\operatorname{Aut}(X,\cdot)$, $\psi\in\operatorname{Aut}(X,\cdot^{\prime})$ such that $\gcd(|\mathcal{G}(X,\cdot)||\mathcal{G}(X,\cdot^{\prime})|,o(\phi)o(\psi))=1$. Then $f:(X,\cdot_{\phi})\to(X,\cdot^{\prime}_{\psi})$ is an isomorphism if and only if $f:(X,\cdot)\to(X,\cdot^{\prime})$ is an isomorphism and $\psi=f\phi f^{-1}$. In particular, $\operatorname{Aut}(X,\cdot_{\phi})$ is precisely the centraliser of $\phi$ in $\operatorname{Aut}(X,\cdot)$.
@@
Assume that $f:(X,\cdot_{\phi})\to(X,\cdot^{\prime}_{\psi})$ is an isomorphism. From 4.6 and the assumptions we find the existence of some $k\in\mathbb{Z}$ such that $(X,(\cdot_{\phi})_{k})=(X,\cdot)$ and $(X,(\cdot^{\prime}_{\psi})_{k})=(X,\cdot^{\prime})$. The functoriality of cabling now yields that $f$ induces an isomorphism $f:(X,\cdot)\to(X,\cdot^{\prime})$. Now for any $x,y\in X$ we find $f(x\cdot_{\phi}y)=f\phi(x\cdot y)$ and $f(x)\cdot_{\psi}f(y)=\psi f(x\cdot y)$, hence $f(x\cdot_{\phi}y)=f(x)\cdot_{\psi}f(y)$ if and only if $\psi=f\phi f^{-1}$. This proves one implication of the statement.
@@ 5. Summary
    $(a,x)\cdot(b,y)=(b+1,y+S\chi_{0}(b)+\Phi(b-a))$
    where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a nonconstant map with $\Phi(0)=0$, $S\in\mathbb{Z}_{p}$ and
    $\chi_{0}:\mathbb{Z}_{p}$ $\to\mathbb{Z}_{p}$
@@
    $(a,x)\cdot(b,y)=(\alpha b+\alpha x,\alpha y+\alpha\Phi(b-a))$
    where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a nonconstant map with $\Phi(x)=\Phi(-x)$ and $\alpha\in\mathbb{Z}_{p}^{\ast}$ is such that $\Phi(\alpha x)=\alpha\Phi(x)$.
@@ 5.1. Indecomposable set-theoretic solutions of size $p^{2}$
Note that these solutions are the same constructed in [6, Theorem 5.1].
@@
Let $p$ be a prime number. Let $t\in\mathbb{Z}_{p}$ be a nonzero element. Let $f:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ be a map such that
@@
Let $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ and $r:X^{2}\to X^{2}$ be the map $r\begin{pmatrix}(i,j)\\ (k,l)\end{pmatrix}=\begin{pmatrix}\lambda_{(i,j)}(k,l)\\ \lambda_{\lambda_{(i,j)}(k,l)}^{-1}(i,j)\end{pmatrix}$, where $\lambda_{(i,j)}(k,l)=(tk+j,t(l-f(tk+j-i)))$. Then $(X,r)$ is a simple solution of the YBE in the following sense: if $(Y,s)$ is a solution and $f:X\twoheadrightarrow Y$ is a surjection with $f(r\times r)=(s\times s)f$, then $|Y|\in\{1,|X|\}$ [^1] This means that the corresponding cycle sets are simple in the sense that there are no nontrivial quotients..
@@
With this notation and fixing $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, it is easy to prove that the map $\Psi(i,j)=(i,-j)$ is an isomorphism of solutions $\Psi:(X,r^{\Phi,\alpha})\to(X,r^{f_{\Phi,\alpha},\alpha^{-1}})$, where $f_{\Phi,\alpha}:i\mapsto-\Phi(\alpha i)$. In fact,
@@ 5.2. Enumeration of indecomposable, irretractable cycle sets of size $p^{2}$
Recall that $\mathcal{F}_{p}$ has been defined as the set of all nonconstant maps $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ such that $\Phi(-A)=\Phi(A)$.
By 5.1, every irretractable cycle sets can be described by a pair $(\Phi,\alpha)$ where $\alpha\in\mathbb{Z}_{p}^{\ast}$ satisfies ${}^{\alpha}\Phi=\Phi$ for all $A\in\mathbb{Z}_{p}$, where ${}^{\alpha}\Phi:A\mapsto\alpha^{-1}\Phi(\alpha A)$. Note that this is an action of $\mathbb{Z}_{p}^{\ast}\cong\mathbb{Z}_{p-1}$ on $\mathcal{F}_{p}$. Furthermore, $(\Phi,\alpha)$ and $(\Phi^{\prime},\alpha^{\prime})$ define isomorphic cycle sets, if and only if $\alpha=\alpha^{\prime}$ and $\Phi^{\prime}={}^{\beta}\Phi$ for some $\beta\in\mathbb{Z}_{p}^{\ast}$.
Assume $p\neq 2$ from now on.
For a pair $(\Phi,\alpha)$ with $\Phi(0)\neq 0$, the condition $\Phi(\alpha A)=\alpha\Phi(A)$ forces $\alpha=1$ and $\Phi$ has to be nonconstant and even. As $\Phi(0)\neq 0$, there is exactly one $\beta\in\mathbb{Z}_{p}^{\ast}$ such that ${}^{\beta}\Phi(0)=1$. Therefore, each cycle set with parameters $(\Phi,\alpha)$, $\Phi(0)\neq 0$, is isomorphic to a unique cycle set with parameters $(\tilde{\Phi},1)$ where $\tilde{\Phi}(0)=1$, and these parameters define mutually nonisomorphic cycle sets. We therefore find $n_{p}=p^{\frac{p-1}{2}}-1$ such cycle sets.
@@
In order to count the number of orbits under the action of $\mathbb{Z}_{p}^{\ast}$, we use Burnside’s lemma. Note that only the elements of $\zeta_{l}$ have fixed points in $\mathcal{F}_{p}$, so we can restrict to those. Given $\alpha\in\zeta_{l}$ with $\operatorname{ord}(\alpha)=d$, a function $\Phi\in\mathcal{F}_{p}$ that satisfies ${}^{\alpha}\Phi=\Phi$, is already defined by its values on coset representatives of $\mathbb{Z}_{p}^{\ast}/\left\langle-1,\alpha\right\rangle$. Under the restriction that $\Phi$ is nonconstant and $\Phi(0)=0$, there are $p^{\frac{p-1}{2d}}-1$ choices for $\Phi$.
As $\left\langle\alpha\right\rangle$ leaves each of these options invariant, we only need to count orbits with respect to the induced action of $\mathbb{Z}_{p}^{\ast}/\left\langle\alpha\right\rangle$ on $\operatorname{Fix}_{\mathcal{F}_{p}}(\alpha)$. Given an element $[\beta]\in\mathbb{Z}_{p}^{\ast}/\left\langle\alpha\right\rangle$, it has fixed points in $\operatorname{Fix}_{\mathcal{F}_{p}}(\alpha)$ if and only if $\beta\in\zeta_{l}$. Assuming the latter, and letting $c=\operatorname{ord}_{\mathbb{Z}_{p}^{\ast}}(\beta)$, we get $p^{\frac{p-1}{2c}}-1$ elements in $\operatorname{Fix}_{\mathcal{F}_{p}}([\beta])$.
@@
equivalence classes for parameters of the form $(\Phi,\alpha)$ where $\operatorname{ord}(\alpha)=d$. Considering that there are $\varphi(d)$ such $\alpha\in\zeta_{l}$, we get the following number of nonisomorphic cycle sets with parameters $(\Phi,\alpha)$, $\Phi(0)\neq 0$:
@@
is a convolution of multiplicative functions [^2] Here, multiplicative means $\psi(mn)=\psi(m)\psi(n)$ for coprime positive integers $m,n$. and therefore also a multiplicative function which evaluates on prime powers $q^{\nu}$, $\nu\geqslant 1$, as
@@
The total number of indecomposable, nonisomorphic, irretractable cycle sets of size $p^{2}$ can therefore be described as:

### v2

@@ 4. Indecomposable irretractable cycle sets of size $p^{2}$ The general case
From the original definition of $\operatorname{Fix}_{A}(L\cap G)$ we see that $\operatorname{Fix}_{A}(L\cap G)^{+}$ is a group. Let $a\in\operatorname{Fix}_{A}(L\cap G)$ and $b\in L\cap G$. Then $\lambda_{\overline{a}}(b)\in L$, but as $\lambda_{\overline{a}}(b)={}^{\overline{a}}b$, also $\lambda_{\overline{a}}(b)\in G$. We conclude that $\lambda_{\overline{a}}(b)\in L\cap G$. Using the alternative description of $\operatorname{Fix}_{A}(L\cap G)$ we now see that $\operatorname{Fix}_{A}(L\cap G)$ is closed under the $\circ$-operation and non-empty. As $A$ is finite, we conclude that $\operatorname{Fix}_{A}(L\cap G)^{\circ}$ is a subgroup of $A^{\circ}$. Therefore $\operatorname{Fix}_{A}(L\cap G)$ is a subbrace. As ${}^{\overline{a}}b\in L\cap G$ for all $a\in\operatorname{Fix}_{A}(L\cap G)$, $b\in L\cap G$ we find indeed that $\operatorname{Fix}_{A}(L\cap G)$ is contained in the normaliser of $L\cap G$ in $A^{\circ}$. ∎
In the remainder of the section, we let $X$ be an indecomposable irretractable cycle set of size $p^{2}$ and $\mathcal{G}=\mathcal{G}(X)$. As we already covered the case where $\mathcal{G}$ is a $p$-group in Section 3, we can assume that $\mathcal{G}$ is not a $p$-group, but recall that $\mathcal{G}$ is solvable by [9, Theorem 2.14]. We associate $X$ with its image in $\mathcal{G}$, which is a transitive cycle base. As $X$ is irretractable, it follows from [2, Lemma 2.1.] that $\operatorname{Soc}(\mathcal{G})=0$.
@@
From now on we consider the unique block system of $X$ under the action of $\mathcal{G}$, and recall that the uniqueness is guaranteed by 1.7. As before we denote this by $\{\mathcal{B}_{x}\mid x\in X\}$. We denote the subgroup of $\mathcal{G}^{\circ}$ that fixes the blocks setwise by $\mathcal{A}$. Also, we define $\mathcal{A}_{p}=\mathcal{A}\cap\mathcal{G}_{p}$ and $\mathcal{A}_{p^{\prime}}=\mathcal{A}\cap\mathcal{G}_{p^{\prime}}$. Note that $\mathcal{A}_{p}$ is a $p$-Sylow subgroup of $\mathcal{A}$ and $\mathcal{A}_{p^{\prime}}$ is a Hall $p^{\prime}$-subgroup of $\mathcal{A}$. In particular, $\mathcal{A}_{p}$ is normal in $\mathcal{G}^{\circ}$ by 1.8.
@@
By 1.8 we find that $|\mathcal{G}_{p^{\prime}}|<p$ and thus the $\lambda$-action of $\mathcal{G}_{p}$ on $\mathcal{G}_{p^{\prime}}$ is trivial. As this same action is transitive on $Z$ we find $|Z|=1$. Let $z\in Z$, then $\lambda_{z}$ is a brace automorphism of $\mathcal{G}_{p}$ by 1.1 and therefore also its restriction to $Y$ yields a cycle set automorphism. For any $x_{i},x_{j}\in X$ we find
@@
We remark that the idea of starting from a finite cycle set $X$ and then considering its projection $Y$ onto the $p$-Sylow subgroup $\mathcal{G}_{p}$, is strongly related to the notion of cabling, see [15, 10]. Recall that for $k\in\mathbb{Z}$, the $k$-cabled cycle set is defined as $(X,\cdot_{k})$ with $x\cdot_{k}y=(k\sigma_{x})(y)$, where $k\sigma_{x}\in\mathcal{G}(X)$ is the $k$-th power of $\sigma_{x}$ in $(\mathcal{G}(X),+)$.
@@
Let $|\mathcal{G}(X,\cdot)|=p^{r}m$ where $(p,m)=1$. As the additive $p$-Sylow subgroup is $\mathcal{G}(X,\cdot)_{p}=\{kg\mid g\in\mathcal{G}\}$ for any non-zero multiple $k$ of $m$, we find that $\mathcal{G}(X,\cdot)_{p}=\mathcal{G}(X,\cdot_{k})$. In particular, $\operatorname{Ret}(X,\cdot_{k})$ isomorphic to the cycle set $\{k\sigma_{x}\mid x\in X\}\subseteq\mathcal{G}(X,\cdot)$ where the equivalence class of $x$ is mapped to $k\sigma_{x}$. If we let $k$ be such that $k\equiv 1\pmod{p^{r}}$ then we find that the $k\sigma_{x}$ is precisely the projection of $\sigma_{x}$ onto $\mathcal{G}(X,\cdot)_{p}$, hence in this case we find that $Y\cong\operatorname{Ret}(X,\cdot_{k})$, with $Y$ as before.
@@
By the functoriality of $G(X,r)$ we get an induced automorphism $\phi^{\prime}$ of $G(X,r)$ which restricts to $\phi$ on the generating set $X\subseteq G(X,\cdot)$. We let $\mathbb{Z}$ denote the trivial brace on $\mathbb{Z}$ and we let $\mathbb{Z}^{\circ}$ act on $G(X,\cdot)$ by $1\cdot g=(\phi^{\prime})^{-1}(g)$ for $g\in G(X,\cdot)$. Hence we can construct the semi-direct product $G(X,\cdot)\rtimes\mathbb{Z}$ in the sense of [20, Corollary 3.37]. The set $X+1\subseteq G(X,\cdot)\rtimes\mathbb{Z}$ is closed under the $\lambda$-action. Hence it is a sub-cycle set which is precisely the cycle set in the statement under the correspondence $x\mapsto x+1$.
@@
Let $(X,\cdot)$ be a finite cycle set and $\phi$ an automorphism of $(X,\cdot)$ of order $m$ coprime to $|\mathcal{G}(X)|$ such that $\phi$ has a fixed point. Then $\mathcal{G}(X,\cdot_{\phi})^{\circ}=\mathcal{G}(X,\cdot)^{\circ}\rtimes\langle\phi\rangle$ as subgroups of $\operatorname{Sym}_{X}$ and $\mathcal{G}(X,\cdot_{\phi})\cong\mathcal{G}(X,\cdot)\rtimes\mathbb{Z}_{m}$ as braces where $\mathbb{Z}_{m}$ is given the trivial brace structure and acts on $\mathcal{G}(X,\cdot)$ by $\phi$.
@@
Let $(X,\cdot)$, $(X,\cdot^{\prime})$ be finite cycle sets and $\phi\in\operatorname{Aut}(X,\cdot)$, $\psi\in\operatorname{Aut}(X,\cdot^{\prime})$ such that $\gcd(|\mathcal{G}(X,\cdot)||\mathcal{G}(X,\cdot^{\prime})|,o(\phi)o(\psi))=1$. Then $f:(X,\cdot_{\phi})\to(X,\cdot^{\prime}_{\psi})$ is an isomorphism if and only if $f:(X,\cdot)\to(X,\cdot^{\prime})$ is an isomorphism and $\psi=f\phi f^{-1}$. In particular, $\operatorname{Aut}(X,\cdot_{\phi})$ is precisely the centraliser of $\phi$ in $\operatorname{Aut}(X,\cdot)$.
@@
Assume that $f:(X,\cdot_{\phi})\to(X,\cdot^{\prime}_{\psi})$ is an isomorphism. From 4.6 and the assumptions we find the existence of some $k\in\mathbb{Z}$ such that $(X,(\cdot_{\phi})_{k})=(X,\cdot)$ and $(X,(\cdot^{\prime}_{\psi})_{k})=(X,\cdot^{\prime})$. The functoriality of cabling now yields that $f$ induces an isomorphism $f:(X,\cdot)\to(X,\cdot^{\prime})$. For any $x,y\in X$ we find $f(x\cdot_{\phi}y)=f\phi(x\cdot y)$ and $f(x)\cdot_{\psi}f(y)=\psi f(x\cdot y)$, hence $f(x\cdot_{\phi}y)=f(x)\cdot_{\psi}f(y)$ if and only if $\psi=f\phi f^{-1}$. This proves one implication of the statement.
@@
Remark 4.9.
Observe that if $\alpha\neq 1$, then $\pi(\mathcal{G}(X))\neq\pi(X)$. Hence the permutation braces of these solutions are all examples of singular brace as defined in [19].
@@ 5. Summary
    $(a,x)\cdot(b,y)=(b+1,y+\chi_{0}(b)S+\Phi(b-a)),$
    where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a non-constant map with $\Phi(0)=0$, $S\in\mathbb{Z}_{p}$ and $\chi_{0}:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ with
@@
    $(a,x)\cdot(b,y)=(\alpha b+\alpha x,\alpha y+\alpha\Phi(b-a)),$
    where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a non-constant map with $\Phi(x)=\Phi(-x)$ and $\alpha\in\mathbb{Z}_{p}^{\ast}$ is such that $\Phi(\alpha x)=\alpha\Phi(x)$.
@@ 5.1. Indecomposable set-theoretic solutions of size $p^{2}$
In fact, these solutions are isomorphic to those constructed in [6, Theorem 5.1] as we will show in the remainder of this section. In particular, this answers [6, Question 7.4] affirmatively.
@@
Let $p$ be a prime number. Let $t\in\mathbb{Z}_{p}$ be a non-zero element. Let $f:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ be a map such that
@@
Let $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ and $r:X^{2}\to X^{2}$ be the map $r\begin{pmatrix}(i,j)\\ (k,l)\end{pmatrix}=\begin{pmatrix}\lambda_{(i,j)}(k,l)\\ \lambda_{\lambda_{(i,j)}(k,l)}^{-1}(i,j)\end{pmatrix}$, where $\lambda_{(i,j)}(k,l)=(tk+j,t(l-f(tk+j-i)))$. Then $(X,r)$ is a simple solution of the YBE in the following sense: if $(Y,s)$ is a solution and $f:(X,r)\twoheadrightarrow(Y,s)$ is a surjective homomorphism, then $|Y|\in\{1,|X|\}$.
@@
With this notation and fixing $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, it is easy to prove that the map $\Psi(i,j)=(i,-j)$ is an isomorphism of solutions $\Psi:(X,r^{\Phi,\alpha})\to(X,r^{f_{\Phi,\alpha},\alpha^{-1}})$, where $f_{\Phi,\alpha}:i\mapsto-\Phi(\alpha i)$, since
@@ 5.2. Enumeration of indecomposable, irretractable cycle sets of size $p^{2}$
Recall that $\mathcal{F}_{p}$ has been defined as the set of all non-constant maps $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ such that $\Phi(-A)=\Phi(A)$.
By 5.1, every irretractable cycle sets can be described by a pair $(\Phi,\alpha)$ where $\alpha\in\mathbb{Z}_{p}^{\ast}$ satisfies ${}^{\alpha}\Phi=\Phi$ for all $A\in\mathbb{Z}_{p}$, where ${}^{\alpha}\Phi:A\mapsto\alpha^{-1}\Phi(\alpha A)$. Note that this is an action of $\mathbb{Z}_{p}^{\ast}\cong\mathbb{Z}_{p-1}$ on $\mathcal{F}_{p}$. Furthermore, $(\Phi,\alpha)$ and $(\Phi^{\prime},\alpha^{\prime})$ define isomorphic cycle sets, if and only if $\alpha=\alpha^{\prime}$ and $\Phi^{\prime}={}^{\beta}\Phi$ for some $\beta\in\mathbb{Z}_{p}^{\ast}$. It follows directly that for $p=2$ we find 2 non-isomorphic indecomposable irretractable cycle sets of size $4$.
Assume $p\neq 2$ from now on. For a pair $(\Phi,\alpha)$ with $\Phi(0)\neq 0$, the condition $\Phi(\alpha A)=\alpha\Phi(A)$ forces $\alpha=1$ and $\Phi$ has to be non-constant and even. As $\Phi(0)\neq 0$, there is exactly one $\beta\in\mathbb{Z}_{p}^{\ast}$ such that ${}^{\beta}\Phi(0)=1$. Therefore, each cycle set with parameters $(\Phi,\alpha)$, $\Phi(0)\neq 0$, is isomorphic to a unique cycle set with parameters $(\tilde{\Phi},1)$ where $\tilde{\Phi}(0)=1$, and these parameters define mutually non-isomorphic cycle sets. We therefore find $n_{p}=p^{\frac{p-1}{2}}-1$ such cycle sets.
@@
In order to count the number of orbits under the action of $\mathbb{Z}_{p}^{\ast}$, we use Burnside’s lemma. Note that only the elements of $\zeta_{l}$ have fixed points in $\mathcal{F}_{p}$, so we can restrict to those. Given $\alpha\in\zeta_{l}$ with $\operatorname{ord}(\alpha)=d$, a function $\Phi\in\mathcal{F}_{p}$ that satisfies ${}^{\alpha}\Phi=\Phi$, is already defined by its values on coset representatives of $\mathbb{Z}_{p}^{\ast}/\left\langle-1,\alpha\right\rangle$. Under the restriction that $\Phi$ is non-constant and $\Phi(0)=0$, there are $p^{\frac{p-1}{2d}}-1$ choices for $\Phi$.
As $\left\langle\alpha\right\rangle$ leaves each of these options invariant, we only need to count orbits with respect to the induced action of $\mathbb{Z}_{p}^{\ast}/\left\langle\alpha\right\rangle$ on $\operatorname{Fix}_{\mathcal{F}_{p}}(\langle\alpha\rangle)$. Given an element $[\beta]\in\mathbb{Z}_{p}^{\ast}/\left\langle\alpha\right\rangle$, it has fixed points in $\operatorname{Fix}_{\mathcal{F}_{p}}(\langle\alpha\rangle)$ if and only if $\beta\in\zeta_{l}$. Assuming the latter, and letting $c=\operatorname{ord}_{\mathbb{Z}_{p}^{\ast}}(\beta)$, we get $p^{\frac{p-1}{2c}}-1$ elements in $\operatorname{Fix}_{\mathcal{F}_{p}}(\langle[\beta]\rangle)$.
@@
equivalence classes for parameters of the form $(\Phi,\alpha)$ where $\operatorname{ord}(\alpha)=d$. Considering that there are $\varphi(d)$ such $\alpha\in\zeta_{l}$, we get the following number of non-isomorphic cycle sets with parameters $(\Phi,\alpha)$, $\Phi(0)\neq 0$:
@@
is a convolution of multiplicative functions. Here, multiplicative means $\mu(mn)=\mu(m)\mu(n)$ for coprime positive integers $m,n$. So also $\psi$ is a multiplicative function which evaluates on prime powers $q^{\nu}$, $\nu\geqslant 1$, as
@@
The total number of indecomposable, non-isomorphic, irretractable cycle sets of size $p^{2}$ can therefore be described as: